OpenStudy (anonymous):

need to simplify -24/36 to get -2/3

OpenStudy (anonymous):

12 goes into 24 twice and 12 goes into 36 three times

OpenStudy (anonymous):

how do I show that though

OpenStudy (anonymous):

try bumping up this question to the top cause 24/12=2 and 36/12=3 to me that is self explanitory lol

OpenStudy (anonymous):

Did both the numerator and the denominator by their largest common factor. The largest common factor of 24 and 36 is 12. So we divided the numerator by 12, and the denominator by 12. Doing so gets us: -24/36 = -2/3

OpenStudy (anonymous):

Ignore my first post. I made a typo in it. Just the post directly above THIS one should be looked at. @C.Hernandez